it takes rob 4 hours to paint a room. it takes darrin 3 hours to paint a room the same size. how long it take them to paint a similar sized room together
Found 2 solutions by stanbon, mananth:
Answer by stanbon(75887)

it takes rob 4 hours to paint a room. it takes darrin 3 hours to paint a room the same size. how long it take them to paint a similar sized room together
-----
Rob rate = 1/4 job/hr
Darrin rate = 1/3 job/hr
Together rate = 1/x job/hr
-------------------------------
Equation:
rate + rate = together rate
1/4 + 1/3 = 1/x
Multiply thru by 12x to get:
3x + 4x = 12
7x = 12
x = 12/7
x = 1hr and (5/12)60 min
---
x = 1 hr and 25 min
================================================
Cheers,
Stan H.

Answer by mananth(16949)

Rob 4 hours
Darrin 3 hours
..
Rob does 1/4 of the job in 1 hour
Darrin does1/3 of the job in 1 hour.
..
both together will do 1/4 +1/3 of the job in 1 hour
..
1/4 +1/3 = 7/12 of the job in 1 hour.
they will take 12 /7 hours to do the job together.
=1.7 hours
...
m.ananth@hotmail.ca